# LanguageBind: Extending Video-Language Pretraining to N-modality by Language-based Semantic Alignment

### v1

@@ Abstract
The video-language (VL) pretraining has achieved remarkable improvement in multiple downstream tasks. However, the current VL pretraining framework is hard to extend to multiple modalities (N modalities, $N\geq 3$) beyond vision and language. We thus propose LanguageBind, taking the language as the bind across different modalities because the language modality is well-explored and contains rich semantics. Specifically, we freeze the language encoder acquired by VL pretraining and then train encoders for other modalities with contrastive learning. As a result, all modalities are mapped to a shared feature space, implementing multi-modal semantic alignment. While LanguageBind ensures that we can extend VL modalities to N modalities, we also need a high-quality dataset with alignment data pairs centered on language. We thus propose VIDAL-10M with 10 Million data with Video, Infrared, Depth, Audio and their corresponding Language. In our VIDAL-10M, all videos are from short video platforms with complete semantics rather than truncated segments from long videos, and all the video, depth, infrared, and audio modalities are aligned to their textual descriptions. After pretraining on VIDAL-10M, we outperform ImageBind by 1.2% R@1 on the MSR-VTT dataset with only 15% of the parameters in the zero-shot video-text retrieval, validating the high quality of our dataset. Beyond this, our LanguageBind has greatly improved in the zero-shot video, audio, depth, and infrared understanding tasks. For instance, on the LLVIP and NYU-D datasets, LanguageBind outperforms ImageBind-huge with 23.8% and 11.1% top-1 accuracy. Code address: https://github.com/PKU-YuanGroup/LanguageBind
@@ 1 Introduction
The proposed LanguageBind ensures that we can extend vision-language to multiple (N) modalities, and our dataset VIDAL-10M benefits more downstream tasks beyond VL tasks, including video retrieval (Luo et al., 2022), depth classification (Cao et al., 2017), infrared classification (Baffa & Lattari, 2018) and audio classification (Palanisamy et al., 2020). In text-to-video zero-shot retrieval on the MSR-VTT dataset, LanguageBind outperforms ImageBind with a 1.2% higher R@1 score, while using less than 15% of the parameters. On the MSVD dataset, LanguageBind achieves a notable R@1 score of 42.7%, surpassing the previous state-of-the-art model, CLIP4Clip (Luo et al., 2022), with an improvement of 4.2%. For zero-shot classification on depth and infrared data, LanguageBind achieves a substantial performance advantage over ImageBind. LanguageBind attains top-1 accuracy of 87.2% and 65.1% on LLVIP and NYU-D, respectively, outperforming ImageBind by 23.8% and 11.1%. For zero-shot audio classification tasks, LanguageBind outperforms ImageBind with a 2.7% higher top-1 accuracy on the ESC50 dataset.
@@ 3.1 Multi-modal Encoders
For the video encoder, we adopted the training framework of CLIP4Clip. The video encoder is initialized from ViT-B/32 and more details are in Appendix B. For other modalities besides language, we employ the 24-layer, 1024-dimensional vision transformer with a patch size of 14. The encoders are initialized from OpenCLIP-large (Ilharco et al., 2021). Depth and infrared are treated as RGB images, which are replicated 3 times in the channel dimension to align with RGB images. Following ImageBind, audio data is transformed into spectrograms with a duration of 10 seconds (128 mel-bins) and we repeat and pad the spectrograms. For example, a 4-second spectrogram would be repeated twice and then padded with zero for an additional 2 seconds. Similarly, we also replicate it 3 times in the channel dimension. If the duration exceeds 10 seconds, we randomly sample three 10-second audio segments, each from the front 1/3, middle 1/3, and back 1/3 of the original audio, and finally stack them together.
@@ 4.2 Video and Audio collection and filtering
During the data collection process, we utilize the aforementioned search terms to retrieve video-text pairs and audio-text pairs from relevant platforms, e.g. YouTube Shorts, and Freesound. Regarding video collection, in order to obtain short videos with high-quality textual descriptions, we implemented a filtering mechanism for the title and hashtags. Video samples with titles containing less than 2 words and without video hashtag labels are excluded from our dataset. Moreover, we removed irrelevant words and hashtags, such as ”youtube”, ”fyp”, ”shorts”, etc. Furthermore, to ensure a complete, consistent, and precise depiction of the event within a single full video, we decide to impose a duration limit of 20 seconds. Shorter videos tend to exhibit better scene coherence and event integrity and are more closely aligned with corresponding hashtags and title descriptions. Ultimately, we obtain a short video dataset that encompasses more specific rather than abstract content. Concerning audio collection, we rank the audio list on the Freesound platform based on its similarity to the search terms. Additionally, we conduct filtering operations similar to those for videos, taking into account factors such as audio ratings, download counts, user comments, tags, and duration. This comprehensive approach allows us to curate and refine the audio and video content more effectively.
@@ 4.3 Modality generation and enhancement
The language modality of VIDAL-10M consists of multi-view texts, including title, hashtags, keyframe captions, video captions, and enhanced captions. The detailed text generation and enhancement pipeline is illustrated in Figure 4. Hashtags in VIDAL-10M are specifically designed to highlight the main subjects and actions depicted in the video. These hashtags serve as key indicators, emphasizing the focal points and dynamic elements of the video. However, hashtags alone may not fully capture the spatial information conveyed by the video frames. To address this limitation, we leverage the image captioning model OFA (Wang et al., 2022) to generate supplementary keyframe captions that enrich the spatial information at the keyframe level. These captions also contain local temporal information related to the video content, which is beneficial for visual-text pretraining. Besides spatial information, temporal information concealed within the video is equally significant, providing crucial insights into the progression and sequencing of events within the video. To further supplement the overall thematic and temporal information of the video, we employ the mPLUG-owl model (Ye et al., 2023) to generate video captions based on the combination of video, title, and hashtags. By leveraging the title and hashtags as accurate video labels, we guide the mPLUG-owl model to generate captions that align with the video theme, reducing potential model bias to a certain extent. Furthermore, to extract valuable information from the generated video captions, we utilize the ChatGPT model to refine and enhance the textual description, thereby greatly improving the quality of the text. By incorporating the above text components, multi-view textual descriptions provide a comprehensive and detailed representation of the video content.
@@ 5 Experiments and Results
5.1 Zero-shot classification in multiple modalities
We compare our model with the recent state-of-the-art multi-modal pretraining models, OpenCLIP (Ilharco et al., 2021) and ImageBind (Girdhar et al., 2023) on multi-modal understanding ability tasks in Table 5.1. For infrared-language zero-shot classification, LanguageBind exhibits a noteworthy 23.8% performance advantage over ImageBind on the LLVIP and outperforms OpenCLIP on all three datasets(LLVIP, FLIR V1, and V2). For depth images, our zero-shot results on NYU-D surpass ImageBind by a substantial margin of 11.1% and outperform OpenCLIP by 19.7%. For audio, we outperform ImageBind by 2.7% on the ESC-50 dataset.
@@ 5.1 Zero-shot classification in multiple modalities
tableInfrared-Language, Depth-Language, and Audio-Language zero-shot classification. We report the top-1 classification accuracy for all datasets. Method Parameter Infrared Depth Audio LLVIP FLIR V1 FLIR V2 NYU-D ESC-50 Random - 50.0 25.0 8.3 10.0 2.75 ImageBind 633M 63.4 - - 54.0 66.9 OpenCLIP 305M 82.2 81.2 42.6 45.4 - LanguageBind 305M 87.2 82.9 48.0 65.1 69.6
5.2 Zero-shot retrieval in Video-Language
We conduct video-text retrieval experiments on two classical datasets, MSR-VTT (Xu et al., 2016) and MSVD (Chen & Dolan, 2011). To ensure a fair comparison, we utilized the CLIP4Clip training framework and provided additional details in Appendix B. Our method outperformed existing techniques obviously, as indicated in Table 2. For the text-to-video task, we achieved a higher R@1 of 1.2% than ImageBind with fewer than 15% parameters on MSR-VTT. Moreover, we surpassed CLIP4Clip (Luo et al., 2022) by 5.0% and CLIP-straight (Portillo-Quintero et al., 2021) by 5.8% on MSR-VTT, and by 4.2% and 5.7% on MSVD, respectively. Furthermore, we achieved state-of-the-art performance in video-to-text retrieval tasks, significantly outperforming CLIP-straight.
Table 2: Zero-shot Video-Text Retrieval Performance on MSR-VTT and MSVD datasets. ”R@n” represents recalls at different points, and ”MR” stands for Median Rank, following the methodology from  (Zhu & Yang, 2020; Luo et al., 2020). We focus on reporting the parameters of the vision encoder. Our experiments are based on 3 million video-text pairs of VIDAL-10M.
Dataset Method Parameter Pt-data R@1normal-$\uparrow$ R@5normal-$\uparrow$ R@10normal-$\uparrow$ MRnormal-$\downarrow$
MSR-VTT SupportSet 78M IN1K, IG65M, HT100M 8.7 23.0 31.1 31.0
HD-VILA 60M HD-VILA-100M 14.6 34.4 44.1 15.0
CLIP-straight 86M WIT400M 31.2 53.7 64.0 4.0
CLIP4Clip 86M WIT400M, HT100M-380k 32.0 57.0 66.9 4.0
(Text-to-Video) HiTeA 37M IN21K, CC15M, WV2.5M 34.4 60.0 69.9 -
ImageBind 633M LAION2B 36.1 - - -
Ours 86M WIT400M, VIDAL-10M 37.3 62.9 71.5 3.0
MSVD SupportSet 78M IN1K, IG65M, HT100M 28.4 60.0 72.9 4.0
(Text-to-Video) Frozen 86M CC3M,WV2.5M 33.7 64.7 76.3 3.0
CLIP-straight 86M WIT400M 37.0 64.1 73.8 3.0
CLIP4Clip 86M WIT400M, HT100M-380k 38.5 66.9 76.8 2.0
Ours 86M WIT400M, VIDAL-10M 42.7 71.2 80.5 2.0
MSR-VTT CLIP-straight 86M WIT400M 27.2 51.7 62.6 5.0
(Video-to-Text) Ours 86M WIT400M, VIDAL-10M 37.1 62.6 71.5 3.0
MSVD CLIP-straight 86M WIT400M 59.9 85.2 90.7 1.0
(Video-to-Text) Ours 86M WIT400M, VIDAL-10M 63.9 88.0 93.1 1.0
6 Ablation Study
6.1 Impact of different text sources
In Table 6.1, we conduct various experiments to explore how different text sources impact language modality. We verify the effectiveness of LanguageBind, trained with text from multiple sources, across various modalities. While some text sources yield good results, we discover that a single text source may not be universally suitable for all downstream tasks and datasets. In terms of video and depth modalities, the ChatGPT enhanced caption proves to be advantageous. For infrared images, the OFA performs best in the LLVIP dataset, while the raw caption achieves the highest accuracy in FLIR v1 and v2. That’s why our VIDAL-10M provides multi-view textual descriptions, allowing for flexibility in selecting an appropriate text source that caters to diverse task requirements.
@@ 6.1 Impact of different text sources
6.2 Scaling the size of dataset
Figure 5: Scaling pretraining data size.
@@ 6.3 Training Loss and Architecture
Following ImageBind, we focus on depth modality, which is visual and spatial modality. We report top-1 accuracy on the NYU-D dataset.
Training epochs
We conduct an experiment in Table 3a to study the effect of training epochs which shows that the LoRA fine-tuning is highly effective. Although 3 epochs of training regimen yield superior accuracy, we chose to optimize for a single epoch, achieving a balance between performance and training cost.
Training batch size
In Table 3b, we evaluate the effect of batch size on representation learning. The experiments have shown that a larger batch size is not necessarily better. In fact, a batch size of 1,024 is the most optimal.
Rank of LoRA
In our investigation, we examined prevalent rank configurations for LoRA, as detailed in Table 3c. We observe that smaller rank values lead to more significant performance improvements, whereas larger rank tends to decrease performance. This trend may be attributed to the potential overfitting of the model.
Temperature for loss
We scrutinized the impact of diverse temperature in Table 3d. We found that the learnable temperature initiated from 0.07 performs best, outperforming the fixed temperature strategy proposed by ImageBind.
Masked ratio
We explore the impact of different mask ratios in Table 3e. The results show that a mask ratio of 0.5 demonstrates the highest performance, requiring only a quarter of the computational resources, aligning with findings in FLIP (Li et al., 2023).
@@
1 3 5
65.1 65.4 64.5
(a) Epochs
512 1k 2k
63.9 65.1 64.5
(b) Batch size
2 4 8
65.1 64.4 64.7
(c) Rank of LoRA
Learn 0.05 0.1
65.1 63.0 62.7
(d) Temperature
0.3 0.5 0.7
64.8 65.1 62.7
(e) Masked ratio

### v7

@@ Abstract
The video-language (VL) pretraining has achieved remarkable improvement in multiple downstream tasks. However, the current VL pretraining framework is hard to extend to multiple modalities (N modalities, $N\geq 3$) beyond vision and language. We thus propose LanguageBind, taking the language as the bind across different modalities because the language modality is well-explored and contains rich semantics. Specifically, we freeze the language encoder acquired by VL pretraining and then train encoders for other modalities with contrastive learning. As a result, all modalities are mapped to a shared feature space, implementing multi-modal semantic alignment. While LanguageBind ensures that we can extend VL modalities to N modalities, we also need a high-quality dataset with alignment data pairs centered on language. We thus propose VIDAL-10M with 10 Million data with Video, Infrared, Depth, Audio and their corresponding Language. In our VIDAL-10M, all videos are from short video platforms with complete semantics rather than truncated segments from long videos, and all the video, depth, infrared, and audio modalities are aligned to their textual descriptions. LanguageBind has achieved superior performance on a wide range of 15 benchmarks covering video, audio, depth, and infrared. Moreover, multiple experiments have provided evidence for the effectiveness of LanguageBind in achieving indirect alignment and complementarity among diverse modalities. Code address: https://github.com/PKU-YuanGroup/LanguageBind
@@ 1 Introduction
Figure 2: LanguageBind achieves superior performances on a broad range of 15 benchmarks across video, audio, depth and infrared.
The proposed LanguageBind ensures that we can extend vision-language to multiple (N) modalities, and our dataset VIDAL-10M benefits more downstream tasks beyond VL tasks, including video retrieval (Luo et al., 2022), depth classification (Cao et al., 2017), infrared classification (Baffa & Lattari, 2018) and audio classification (Palanisamy et al., 2020). As shown in Figure 2, LanguageBind achieves superior performances on a broad range of 15 tasks. In zero-shot text to video retrieval, LanguageBind achieves superior performance on four datasets, surpassing InterVideo (Wang et al., 2022c) by 1.9% on MSR-VTT (Xu et al., 2016), 8.8% on MSVD (Chen & Dolan, 2011), 6.3% on DiDeMo (Anne Hendricks et al., 2017), and 4.4% on ActivityNet (Caba Heilbron et al., 2015). For zero-shot classification on depth and infrared data, LanguageBind achieves a substantial performance advantage over ImageBind. LanguageBind attains top-1 accuracy of 87.2% and 65.1% on LLVIP and NYU-D, respectively, outperforming ImageBind by 23.8% and 11.1%. For zero-shot audio classification tasks, LanguageBind outperforms ImageBind with a 23.8% higher top-1 accuracy on the ESC50 dataset.
@@ 3.1 Multi-modal Encoders
For other modalities besides language, we employ the 24-layer, 1024-dimensional vision transformer with a patch size of 14. The encoders are initialized from OpenCLIP-large (Ilharco et al., 2021). Depth and infrared are treated as RGB images, which are replicated 3 times in the channel dimension to align with RGB images. Following ImageBind, audio data is transformed into spectrograms with a duration of 10 seconds (128 mel-bins) and we repeat and pad the spectrograms. For example, a 4-second spectrogram would be repeated twice and then padded with zero for an additional 2 seconds. Similarly, we also replicate it 3 times in the channel dimension. If the duration exceeds 10 seconds, we randomly sample three 10-second audio segments, each from the front 1/3, middle 1/3, and back 1/3 of the original audio, and finally stack them together.
@@ 4.2 Video and Audio collection and filtering
During the data collection process, we utilize the aforementioned search terms to retrieve video-text pairs and audio-text pairs from relevant platforms, e.g. YouTube Shorts, and Freesound. Regarding video collection, in order to obtain short videos with high-quality textual descriptions, we implemented a filtering mechanism for the title and hashtags. Video samples with titles containing less than 2 words and without video hashtag labels are excluded from our dataset. Moreover, we removed irrelevant words and hashtags, such as ”youtube”, ”fyp”, ”shorts”, etc. Furthermore, to ensure a complete, consistent, and precise depiction of the event within a single full video, we decide to impose a duration limit of 20 seconds. Shorter videos tend to exhibit better scene coherence and event integrity and are more closely aligned with corresponding hashtags and title descriptions. Ultimately, we obtain a short video dataset that encompasses more specific rather than abstract content. Concerning audio collection, we rank the audio list on different audio platforms based on its similarity to the search terms. Additionally, we conduct filtering operations similar to videos, taking into account factors such as audio ratings, download counts, user comments, tags, and duration. This comprehensive approach allows us to curate and refine the audio and video content more effectively.
4.3 Modality generation and enhancement
@@
The language modality of VIDAL-10M consists of multi-view texts, including title, hashtags, keyframe captions, video captions, and enhanced captions. The detailed text generation and enhancement pipeline is illustrated in Figure 7. Hashtags in VIDAL-10M are specifically designed to highlight the main subjects and actions depicted in the video. These hashtags serve as key indicators, emphasizing the focal points and dynamic elements of the video. However, hashtags alone may not fully capture the spatial information conveyed by the video frames. To address this limitation, we leverage the image captioning model OFA (Wang et al., 2022b) to generate supplementary keyframe captions that enrich the spatial information at the keyframe level. These captions also contain local temporal information related to the video content, which is beneficial for visual-text pretraining. Besides spatial information, temporal information concealed within the video is equally significant, providing crucial insights into the progression and sequencing of events within the video. To further supplement the overall thematic and temporal information of the video, we employ the mPLUG-owl model (Ye et al., 2023) to generate video captions based on the combination of video, title, and hashtags. By leveraging the title and hashtags as accurate video labels, we guide the mPLUG-owl model to generate captions that align with the video theme, reducing potential model bias to a certain extent. Furthermore, to extract valuable information from the generated video captions, we utilize the ChatGPT model to refine and enhance the textual description, thereby greatly improving the quality of the text. By incorporating the above text components, multi-view textual descriptions provide a comprehensive and detailed representation of the video content.
@@ 5 Experiments and Results
5.1 Zero-shot retrieval in Video-Language
Comparison to prior methods
In the zero-shot video-text retrieval benchmark, we utilize ViT-L/14 as the video encoder and add temporal attention layers for fair comparison, which can be found in Appendix B. According to the results presented in Table 2, the performance of LanguageBind exceeds that of VideoCoca (Yan et al., 2022) and OmniVL (Wang et al., 2022a) by 8.3% and 8.0% respectively on MSR-VTT. In comparison to the ImageBind model utilizing the Vit-Huge architecture, the LanguageBind model, employing the Vit-Large model, showcases superior experimental outcomes. Furthermore, compared to models based on CLIP-Large but using more training data, LanguageBind achieves state-of-the-art (SOTA) performance on four datasets, outperforming InterVideo (Wang et al., 2022c) by 1.9% on MSR-VTT, 8.8% on MSVD, 6.3% on DiDeMo, and 4.4% on ActivityNet. We also exceed TVTSv2 (Zeng et al., 2023) by 4.4% and 3.2% on MSR-VTT and DiDeMo, respectively. Moreover, we outperforms UMT-L Li et al. (2023a) on all datasets. For a fair comparison of dataset validity, we use the Vit-B/32 model of CLIP4CLIP to conduct validation experiments using the 100K subset of VIDAL-10M and the 380k subset of HowTo100M. As shown in Table 3, the VIDAL-100k outperforms the HT100M-380k on both MSRVTT and MSVD datasets, validating the effectiveness of our dataset.
Zero-shot X-Language classification
We compare our model with the recent state-of-the-art multi-modal pretraining models, OpenCLIP (Ilharco et al., 2021) and ImageBind (Girdhar et al., 2023) on multi-modal understanding ability tasks in Table 5.1. For video zero-shot classification, we outperform ImageBind by 14.0% with a smaller model on Kinetics-400 (Kay et al., 2017), and we also report the results of multi-view/crop (Simonyan & Zisserman, 2014) on OpenCLIP for further comparison. For infrared, LanguageBind exhibits a noteworthy 23.8% performance advantage over ImageBind on the LLVIP and outperforms OpenCLIP on all three datasets (LLVIP, FLIR V1, and V2). For depth images, our zero-shot results on NYU-D surpass ImageBind by a substantial margin of 11.1% and outperform OpenCLIP by 19.7%. For audio, we outperform ImageBind by 10.1% on Audioset  dataset and 1.1% on VGGSound dataset. We outperform ImageBind by a large margin of 23.9% on the ESC-50 dataset.
Table 2: Zero-shot Video-Text retrieval performance of LanguageBind across four datasets. $*{}^{*}$ donates the result of full tuning. ${}^{\dagger}$ donates the results are training on Huge model. 10M donates training with 10M video-text pairs.
Method Dataset MSR-VTT MSVD DiDeMo ActivityNet
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
Non-CLIP models
OmniVL 14M 34.6 58.4 66.6 - - - 33.3 58.7 68.5 - - -
VideoCoCa 100M 34.3 57.8 67.0 - - - - - - 34.5 63.2 76.6
CLIP-H/14
ImageBind - 36.8 61.8 70.0 - - - - - - - - -
CLIP-L/14
UMT 5M 33.3 58.1 66.7 44.4 73.3 82.4 34.0 60.4 68.7 31.9 69.2 72.0
TVTSv2 8.5M 38.2 62.4 73.2 - - - 34.6 61.9 71.5 - - -
InternVideo 12.8M 40.7 - - 43.4 - - 31.5 - - 30.7 - -
LanguageBind 3M 42.6 65.4 75.5 52.2 79.4 87.3 37.8 63.2 73.4 35.1 63.4 76.6
$*{}^{*}$ 3M 42.7 67.1 77.0 53.5 80.5 87.5 38.1 65.0 73.6 36.9 65.1 77.2
$*{}^{*}$ 10M 42.8 67.5 76.0 54.1 81.1 88.1 39.7 65.5 73.8 38.4 66.6 77.9
CLIP-H/14
${}^{*{\dagger}}$ 10M 44.8 70.0 78.7 53.9 80.4 87.8 39.9 66.1 74.6 41.0 68.4 80.0
Table 3: Zero-shot Video-Text retrieval of CLIP4Clip to verify the effectiveness of VIDAL-10M.
Dataset Method Parameter Source R@1normal-$\uparrow$ R@5normal-$\uparrow$ R@10normal-$\uparrow$ MRnormal-$\downarrow$
MSR-VTT CLIP4Clip 86M WIT400M, HT100M-380k 32.0 57.0 66.9 4.0
CLIP4Clip 86M WIT400M, VIDAL-100k 35.7 60.8 71.5 3.0
MSVD CLIP4Clip 86M WIT400M, HT100M-380k 38.5 66.9 76.8 2.0
CLIP4Clip 86M WIT400M, VIDAL-100k 42.0 70.0 79.2 2.0
@@ 5.1 Zero-shot retrieval in Video-Language
tableZero-shot X-Language classification. We report the mAP for Audioset Audio-only (AS-A) (Gemmeke et al., 2017) and top-1 accuracy results for others. $*{}^{*}$ donates the results of full tuning. Method Size Video Infrared Depth Audio K400 K600 LLVIP FLIR V1 FLIR V2 NYU-D AS-A ESC-50 VGGS ImageBind Huge 50.0 - 63.4 - - 54.0 17.6 66.9 27.8 OpenCLIP Large 60.7 59.0 82.2 81.2 42.6 45.4 - - - LanguageBind Large 64.0 61.9 87.2 82.9 48.0 65.1 27.7 91.8 28.9 $*{}^{*}$ Large - - - - - - 30.0 94.0 38.6
5.2 Zero-shot in multiple modalities
Zero-shot Audio-Language retrieval
Table 4: Zero-shot Audio-Language retrieval. $*{}^{*}$ donates the results of full tuning.
Method Clotho Audiocaps
R@1 R@10 R@1 R@10
AVFIC 3.0 17.5 8.7 37.7
ImageBind 6.0 28.4 9.3 42.3
VALOR 8.4 - - -
LanguageBind 12.1 44.0 12.2 53.2
$*{}^{*}$ 16.7 52.0 19.7 67.6
We compare zero-shot text-to-audio retrieval performance on Clotho (Font et al., 2013) datasets and Audiocaps (Kim et al., 2019) dataset in Table 4. For the Clotho dataset, LanguageBind exhibits a significantly higher margin surpassing AVFIC (Nagrani et al., 2022) and ImageBind by 9.1% and 6.1% respectively. Moreover, our LanguageBind model also surpasses the powerful baseline of VALOR (Chen et al., 2023a) by 3.7%. The same trend is also observed in the Audiocaps dataset. LanguageBind outperformed AVFIC and ImageBind by 2.9% and 5.5%, respectively. Overall, LanguageBind significantly outperforms prior works on two benchmarks validating that it is an efficient way to align audio and language modalities.
Zeor-shot langauge-based multi-modal joint retrieval
In Table 5.2, we conduct multi-modal joint retrieval to explore the complementarity of joint space. We report the R@1 scores on MSR-VTT and Place datasets, while reporting accuracy on other datasets. For MSR-VTT, we only evaluate using videos that include audio. Integrating audio embeddings for video-language retrieval further improves performance, increasing it from 41.4 to 42.0. Similar trends have been observed in other modalities, where each modality has the potential to enhance the performance when combined with other modalities. These results demonstrate that LanguageBind is capable of learning a more consistent feature space.
Emergent zero-shot retrieval
As shown in Table 5.2, we explore the zero-shot performance of emergency coverage in four datasets, including RGB images, audio, infrared, and depth. Due to the novelty of our approach, there are no ”fair” baseline models for comparison. Nonetheless, we compare our results with ImageBind, which aligns with images directly. For example, we achieved R@1 scores of 10.6 and 10.0 on AVE (Tian et al., 2018) and VGGS, respectively. On each benchmark, the performance of emergency zero-shot retrieval achieves significant gains, even approaching results obtained by incorporating textual features. These results suggest that LanguageBind aligns various modalities and implicitly transfers text supervision associated with specific modalities and tasks.
Table 5: Comparison of multi-modal language based retrieval. $*{}^{*}$ donates that it is not clear whether only videos with audio are included. ${}^{\dagger}$ donates that dark nighttime images.
Dataset Method Task Top-1
@@ 5.2 Zero-shot in multiple modalities
[…]
Table 6: Comparison of emergent zero-shot retrieval. ${}^{\dagger}$ donates that we randomly select 10% data to test.
Dataset Method Task Emergent R@1
[…]
ImageBind 36.9
[…]
ImageBind 28.7
[…]
@@ 5.3 Training Loss and Architecture
Dataset 1 5
NYU-D 65.1 64.5
LLVIP 83.9 81.1
FLIR V1 82.9 85.0
FLIR V2 48.0 44.7
(a) Training epochs
Dataset 512 1k 2k
NYU-D 63.9 65.1 64.5
LLVIP 80.0 83.9 78.6
FLIR V1 81.6 82.9 85.2
FLIR V2 45.1 48.0 47.9
(b) Training batch size
Scratch Full tuning LoRA
Time 1.4h 1.4h 0.8h
Mems 278M 278M 132M
LLVIP 57.1 85.1 84.8
FLIR V1 74.7 81.3 81.6
FLIR V2 54.4 41.9 46.6
ESC-50 86.8 88.9 87.4
Clotho 8.8 9.8 10.1
(c) Training strategy
Dataset 2 4 8
NYU-D 65.1 64.4 64.7
LLVIP 83.9 78.0 -
FLIR V1 82.9 74.4 -
FLIR V2 48.0 45.8 -
(d) Rank of LoRA
Dataset Learn 0.05 0.1
NYU-D 65.1 63.0 62.7
LLVIP 83.9 81.8 83.1
FLIR V1 82.9 83.3 80.3
FLIR V2 48.0 45.0 43.2
(e) Temperature for loss
Dataset 0.0 0.3 0.5 0.7
NYU-D - 64.8 65.1 62.7
LLVIP 80.3 79.9 83.9 81.5
FLIR V1 83.5 84.2 82.9 81.9
FLIR V2 43.2 44.0 48.0 42.5
(f) Masked ratio
Following ImageBind, we mainly focus on depth and infrared, which are visual and spatial modality. We report R@1 score for Clotho and Audiocaps datasets and top-1 accuracy for others. We provide more ablation in Appendix E
Training epochs. We conduct an experiment in Table 7a to study the effect of training epochs which shows that the LoRA fine-tuning is highly effective. Although 3 epochs of training regimen yield superior accuracy, we chose to optimize for a single epoch, achieving a balance between performance and training cost.
Training batch size. In Table 7b, we evaluate the effect of batch size on representation learning. The experiments have shown that a larger batch size is not necessarily better. In fact, a batch size of 1,024 is the most optimal.
Training strategy. As indicated by Table 7c, we compare three different strategies. Training from scratch exhibits the poorest performance, likely due to the lack of prior knowledge from CLIP pretraining. On the other hand, full tuning shows significant improvement compared to training from scratch. This highlights the positive impact of leveraging prior knowledge in the form of pre-trained weights. Meanwhile, the LoRA method stands out for its advantages in terms of time and memory cost. It requires less time and memory resources compared to full tuning. Furthermore, LoRA outperforms full tuning on multiple datasets such as LLVIP, FLIRv1, and Clotho. This indicates that LoRA is not only efficient but also effective in learning new knowledge specific to different domains while better retaining the previously acquired knowledge from the pre-trained OpenCLIP models.
Rank of LoRA. In our investigation, we examined prevalent rank configurations for LoRA, as detailed in Table 7d. We observe that smaller rank values lead to more significant performance improvements, whereas larger rank tends to decrease performance. This trend may be attributed to the potential overfitting of the model.
Temperature for loss. We scrutinize the impact of diverse temperature in Table 7e. We find that the learnable temperature initiated from 0.07 performs best, outperforming the fixed temperature strategy proposed by ImageBind.
Masked ratio. We explore the impact of different mask ratios in Table 7f. The results show that a mask ratio of 0.5 demonstrates the highest performance, requiring only a quarter of the computational resources, aligning with findings in FLIP (Li et al., 2023b).